OpenStudy (anonymous):

Solve. 6/x=7/x-1/2 The solution is __ Or it has no solution

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

x = 2

OpenStudy (anonymous):

this can be done lke this 6/x-7/x=-1/2 -1/x=-1/2 hence x=2 i hope its correct